# Privacy Beyond Pixels: Latent Anonymization for Privacy-Preserving Video Understanding

### v1

@@ Appendix B Implementation Details
All of our code is implemented in PyTorch paszke2019pytorch. In this section, we provide implementation details regarding network architecture, input preprocessing, hyperparameters, and training schedules.
B.1 Network Architecture
@@
Each video encoder $f_{E}$ model is left unchanged from the original implementation. The $f_{T_{AR}}$ classifier head is a simple linear layer Linear($d$, NN), where $d$ is the feature vector dimension of $f_{E}$ and $N$ is the number of classes in $\mathbb{D}_{reco}$. In the temporal action detection and anomaly detection task, classifier implementations are unmodified from the original TriDet shi2023tridet and MGFN chen2023mgfn works, respectively. For the private attribute prediction task, a 2-layer MLP is used: Linear($d$, $d$) $\rightarrow$ Linear($d$, 7) with a ReLU activation after the first layer. For the $f_{A}$ AAM, we ablate different architecture styles (see Table 8). To break it down, we tried standard MLPs of different depths and self-attention based adapters of different depths. Each MLP layer is composed of a Linear($d$, $d$) followed by a ReLU activation and a BatchNorm1D layer, and dropout with a probability of 0.1 during training. The self-attention layers are standard MulitheadAttention blocks with dim $d$ and 8 heads by default, with the head count ablated in Table 10.
B.2 Inputs and Augmentations
@@
All inputs consist of 16 frame clips sampled with consecutive frames, resized to spatial resolution of $224\times 224$. For training, only random resized crop and random horizontal flip with probability 50% are utilized. In validation, the short edge is resized to 256, then a center crop of $224\times 224$ is taken. Standard ImageNet krizhevsky2012imagenet mean and standard deviation based normalization is applied in both settings. The input and augmentation protocol is consistent for every $f_{E}$.
B.3 Training Details and Hyperparameters
@@
Each AAM variation is trained using an $\ell_{1}$ loss to reconstruct the input features for 100 epochs with the AdamW loshchilov2017decoupled optimizer and a learning rate of 2e-5. Kinetics400 carreira2017quo features are used as the train-test set. Privacy evaluation is carried out using supervised training of the predictor MLP for 100 epochs at a learning rate of 1e-3. A learning rate scheduler is based on the loss plateau where it decreases the learning rate to 1/5th.
@@
For anonymization training, the base learning is 1e-4 for both $f_{A}$ and $f_{T^{*}}$, corresponding to a batch size of 512, scaled when necessary according to the linear scaling rule goyal2017accurate. By default, $\omega_{LC}=100$, $\omega_{T}=1$, and $\omega_{B}=1$ (Main Equation 7). The anonymization training is carried out for 100 epochs.
B.4 Additional Collaborative Task Losses
@@
Temporal Action Detection Loss $\mathcal{L}_{TAD}$ (TriDet shi2023tridet):
@@
where $N_{pos}$ and $N_{neg}$ are the numbers of positive and negative samples, respectively; $\mathds{1}{{c^{l}_{t}>0}}$ is an indicator function that equals 1 if $c^{l}_{t}>0$ (positive sample) and 0 otherwise; $\sigma_{IoU}$ is the temporal Intersection over Union (IoU) between the predicted segment and the ground truth, serving as a weighting factor; $\mathcal{L}_{cls}$ is the classification loss, implemented as the focal loss ross2017focal; and $\mathcal{L}_{reg}$ is the regression loss, implemented as the IoU loss rezatofighi2019generalizediou. The weighting factor $\sigma_{IoU}$ emphasizes predictions with higher temporal IoU, ensuring that higher-quality predictions contribute more significantly during training. Positive samples are determined using center sampling, where instants near the center of an action instance are labeled as positive, and others are considered negative.
Anomaly Detection Loss $\mathcal{L}_{AD}$ (MGFN chen2023mgfn):
@@
with $y$ as the video-level label ($y=1$ for anomaly, $y=0$ for normal) and $s^{i,j}$ as the computed anomaly score for frame $i$ in segment $j$. Following sultani2018real, it incorporate a temporal smoothness term $\mathcal{L}_{ts}$ and a sparsity term $\mathcal{L}_{sp}$:
@@
where $B$ is the batch size, $M$ denotes the feature magnitude of the corresponding segment, $l$ is an indicator function, and $D(\cdot,\cdot)$ is a distance function. Refer to chen2023mgfn for more details.
@@ Appendix C Additional Experiments
Num MHSA
Heads
VISPR HMDB51 UCF Crime
@@
Figure 3 follows singh2022decouple to plot privacy-utility curve (NHV=0.6833) using PA-HMDB test set. Varying task weights leads to controllable trade-off curve.
@@ C.1 Comparison with Other Prior Methods
Previous work wu2020privacy; dave2022spact; fioresi2023ted has already shown that the learnable anonymization techniques outperform methods such as downsampling, blurring, and blackening. Main Table 1 shows a comparison to these techniques using the I3D carreira2017quo model. In Downsample-2x and Downsample-4x, the input frames have their resolution reduced by a factor of 2 ($112\times 112$) and 4 ($56\times 56$). In Blackening and Blurring, subjects are detected using an object detector to detect human subjects and obfuscated using the same methods as described in wu2020privacy; dave2022spact; fioresi2023ted. We see that none of these techniques achieve an acceptable level of anonymization, and almost all reduce utility more than our SPLAVU method, further demonstrating the capability of our framework.
Table 11 shows the results of our proposed method on PA-HMDB wu2020privacy compared to the baseline model on raw data.
@@ C.2 Additional Experiments with Large Foundation Models
Due to the low compute cost and focus on maintaining the capabilities of powerful models, our SPLAVU framework is able to scale up to the largest video foundational models currently available. Table 12 demonstrates the high privacy-utility tradeoff achieved by our method using InternVideo-H wang2022internvideo, and Main Paper Table 1 shows results using VideoMAEv2-G wang2023videomaeV2. In these experiments, action recognition performance was exactly maintained, and private attribute prediction was dropped more than for the smaller models, with only a modest reduction in temporal action detection performance.
@@
Anonymization
Method
Model VISPR HMDB51 UCF101 THUMOS14
@@ C.2.1 Training Compute
One of the many benefits of our SPLAVU framework is its very low compute/training cost. Table 13 shows the overall count of trainable parameters for previous frameworks compared to our AAM. For VideoMAE-Base, our SPLAVU framework with the self-attention AAM has 88.7% less trainable parameters when compared to existing approaches. This difference is even greater when scaling to larger models. Training less parameters can reduce the tendency to overfit on the proxy task and allow for learning an effective anonymization on limited training data (see Main Paper Table 3). Also, in federated learning, these parameters are communicated between the server and clients, so the reduced learnable parameters are useful in efficient and privacy-preserving federated learning zhao2023fedprompt; yu2022differentially.
@@ C.2.2 Precomputing Feature Embeddings
Training
Data
PAP AR TAD AD Training
@@ Appendix D Training Algorithm
Algorithm 1 formalizes the SPLAVU workflow notation. We consider anonymizer $f_{A}$ and task heads $f_{T_{AR}}$, $f_{T_{TAD}}$, and $f_{T_{AD}}$ for the anonymization training and $f_{reco}$, $f_{tad}$, and $f_{wsad}$ for downstream tasks. In order, these models are parameterized by $\theta_{A}$, $\theta_{{T}_{AR}}$, $\theta_{{T}_{TAD}}$, $\theta_{{T}_{AD}}$, $\theta_{reco}$, $\theta_{tad}$, and $\theta_{wsad}$. $\mathbb{D}_{reco}$, $\mathbb{D}_{tad}$, and $\mathbb{D}_{wsad}$ are all used in the proxy anonymization process, then also for the downstream task evaluation. The downstream $\mathbb{D}_{reco}$ may be the same or different from during the anonymization process. $\mathbb{D}_{anomaly}$ and $\mathbb{D}_{tad}$ are the anomaly detection and temporal action detection datasets used in their respective task evaluation.
1Anonymization Training
2
3Inputs:
4Datasets: $\mathbb{D}_{reco}$, $\mathbb{D}_{tad}$, $\mathbb{D}_{wsad}$
5# of Epochs: anon_epochsanon\_epochs
6Learning Rates: $\alpha_{A}$, $\alpha_{AR}$, $\alpha_{TAD}$, $\alpha_{AD}$
7Hyperparameters: $\omega_{A}$, $\omega_{T}$, $\omega_{B}$, $\omega_{LC}$, $\omega_{AR}$, $\omega_{TAD}$, $\omega_{AD}$
8Output: $\theta_{A}$, $\theta_{{T}_{AR}}$, $\theta_{{T}_{TAD}}$, $\theta_{{T}_{AD}}$
9Model Initialization:
10Initialize $f_{E}$ with Kinetics400 weights carreira2017quo;
11Initialize $\theta_{A}\leftarrow\theta_{A}-\alpha_{A}\nabla_{\theta_{A}}(\mathcal{L}_{L1}(\theta_{A}))$
12Multitask Anonymization Training:
13for $e_{0}\leftarrow 1$ to anon_epochsanon\_epochs do
14 $\theta_{A}\leftarrow\theta_{A}-\alpha_{A}\nabla_{\theta_{A}}(\omega_{LC}\mathcal{L}_{LC}(\theta_{A})+\omega_{T}\mathcal{L}_{T^{*}}(\theta_{A},\theta_{T_{AR}},\theta_{T_{TAD}},\theta_{T_{AD}})-\omega_{B}L_{B}(\theta_{A}))$ $\theta_{T_{AR}}\leftarrow\theta_{T_{AR}}-\alpha_{AR}\nabla_{\theta_{T_{AR}}}(\mathcal{L}_{AR}(\theta_{T_{AR}},\theta_{A})),$ $\theta_{T_{TAD}}\leftarrow\theta_{T_{TAD}}-\alpha_{TAD}\nabla_{\theta_{T_{TAD}}}(\mathcal{L}_{TAD}(\theta_{T_{TAD}},\theta_{A})),$ $\theta_{T_{AD}}\leftarrow\theta_{T_{AD}}-\alpha_{AD}\nabla_{\theta_{T_{AD}}}(\mathcal{L}_{AD}(\theta_{T_{AD}},\theta_{A})),$
15 end for
16
17 Downstream Tasks Evaluation
18
19
20Inputs:
21Datasets: $\mathbb{D}_{reco}$, $\mathbb{D}_{anomaly}$, $\mathbb{D}_{tad}$
22# of Epochs: reco_epochsreco\_epochs, anomaly_epochsanomaly\_epochs, tad_epochstad\_epochs
23Learning Rates: $\alpha_{reco}$, $\alpha_{wsad}$, $\alpha_{tad}$
24
25Output: $\theta_{reco}$, $\theta_{wsad}$, $\theta_{tad}$
26 Privacy-Preserved Action Recognition Training:
27for $e_{0}\leftarrow 1$ to reco_epochsreco\_epochs do
28 $\theta_{reco}\leftarrow\theta_{reco}-\alpha_{reco}\nabla_{\theta_{reco}}(\mathcal{L}_{T}(\theta_{reco},\theta_{A})),$
29 end for
30
31 Feature Extraction on $\mathbb{D}_{anomaly}$:
[…]
@@ 31 Feature Extraction on $\mathbb{D}_{anomaly}$:
33 Privacy-Preserved Weakly-Supervised Anomaly Detection (WSAD) Training:
34for $e_{0}\leftarrow 1$ to anomaly_epochsanomaly\_epochs do
35 $\theta_{wsad}\leftarrow\theta_{wsad}-\alpha_{wsad}\nabla_{\theta_{wsad}}(L_{wsad}(\theta_{wsad},\mathbb{F}_{anomaly}))$
36 end for
37
38 Feature Extraction on $\mathbb{D}_{tad}$:
[…]
40 Privacy-Preserved Temporal Action Detection (TAD) Training:
41for $e_{0}\leftarrow 1$ to tad_epochstad\_epochs do
42 $\theta_{tad}\leftarrow\theta_{tad}-\alpha_{tad}\nabla_{\theta_{tad}}(L_{tad}(\theta_{tad},\mathbb{F}_{anomaly}))$
43 end for
44

### v2

@@ Appendix B Implementation Details
All of our code is implemented in PyTorch Paszke et al. (2019). In this section, we clarify notation and provide implementation details regarding network architecture, input preprocessing, hyperparameters, and training schedules.
B.1 Clarifying Global Embedding Notation
In the main paper, we use the notation $\mathbf{h}_{t}=f_{E}(\mathbf{x}_{t})$ to denote the global clip-level embedding produced by the frozen video encoder $f_{E}$. In standard practice, transformer-based video encoders (e.g., VideoMAE, ViViT) produce a sequence of tokens prepended by a special [CLS] token: $f_{E}(\mathbf{x}_{t})=[\mathbf{h}_{t}^{CLS};\mathbf{h}_{t}^{0};\ldots;\mathbf{h}_{t}^{N}]$, where $\mathbf{h}_{t}^{\text{CLS}}$ is the global classification token and $\mathbf{h}_{t}^{0},\ldots,\mathbf{h}_{t}^{N}$ denote patchwise spatio-temporal tokens. In CNN-based video encoders (e.g., I3D, R3D), the encoder processes the input clip through a hierarchy of spatiotemporal convolutional blocks, producing a feature tensor $f_{E}(\mathbf{x}_{t})=\mathbf{F}_{t}\in\mathbb{R}^{T^{\prime}\times H^{\prime}\times W^{\prime}\times C},$ which is then aggregated into a global clip-level representation via spatiotemporal average pooling: $\mathbf{h}_{t}^{\text{pool}}=\text{AvgPool}(\mathbf{F}_{t})\in\mathbb{R}^{C}.$ This pooled feature $\mathbf{h}_{t}^{\text{pool}}$ serves as the analog to the [CLS] token in transformer-based encoders. Because our anonymizer operates on the unified notion of a single clip-level representation, either the [CLS] token or the globally pooled CNN feature, we simplify notation by writing $f_{E}(\mathbf{x}_{t})$ to refer to this global embedding regardless of encoder type.
B.2 Network Architecture
@@
Each video encoder $f_{E}$ model is left unchanged from the original implementation. The $f_{T_{AR}}$ classifier head is a simple linear layer Linear($d$, NN), where $d$ is the feature vector dimension of $f_{E}$ and $N$ is the number of classes in $\mathbb{D}_{AR}$. In the temporal action detection and anomaly detection task, classifier implementations are unmodified from the original TriDet Shi et al. (2023) and MGFN Chen et al. (2023) works, respectively. For the private attribute prediction task, a 2-layer MLP is used: Linear($d$, $d$) $\rightarrow$ Linear($d$, 7) with a ReLU activation after the first layer. For the $f_{A}$ AAM, we ablate different architecture styles (see Table 10). To break it down, we tried standard MLPs of different depths and self-attention based adapters of different depths. Each MLP layer is composed of a Linear($d$, $d$) followed by a ReLU activation and a BatchNorm1D layer, and dropout with a probability of 0.1 during training. The self-attention layers are standard MultiheadAttention blocks with dim $d$ and 8 heads by default, with the head count ablated in Table 12.
B.3 Inputs and Augmentations
@@
All inputs consist of 16 frame clips sampled with consecutive frames, resized to spatial resolution of $224\times 224$. For training, only random resized crop and random horizontal flip with probability 50% are utilized. In validation, the short edge is resized to 256, then a center crop of $224\times 224$ is taken. Standard ImageNet Krizhevsky et al. (2012) mean and standard deviation based normalization is applied in both settings. The input and augmentation protocol is consistent for every $f_{E}$.
B.4 Training Details and Hyperparameters
@@
Each AAM variation is trained using an $\ell_{1}$ loss to reconstruct the input features for 100 epochs with the AdamW Loshchilov and Hutter (2017) optimizer and a learning rate of 2e-5. Kinetics400 Carreira and Zisserman (2017) features are used as the train-test set. Privacy evaluation is carried out using supervised training of the predictor MLP for 100 epochs at a learning rate of 1e-3. A learning rate scheduler is based on the loss plateau where it decreases the learning rate to 1/5th.
@@
For anonymization training, the base learning is 1e-4 for both $f_{A}$ and $f_{T^{*}}$, corresponding to a batch size of 512, scaled when necessary according to the linear scaling rule Goyal et al. (2017a). By default, $\omega_{LC}=100$, $\omega_{T}=1$, and $\omega_{B}=1$ (Main Equation 7). The anonymization training is carried out for 100 epochs.
B.5 Additional Collaborative Task Losses
@@
Temporal Action Detection Loss $\mathcal{L}_{TAD}$ (TriDet Shi et al. (2023)):
@@
where $N_{pos}$ and $N_{neg}$ are the numbers of positive and negative samples, respectively; $\mathds{1}{{c^{l}_{t}>0}}$ is an indicator function that equals 1 if $c^{l}_{t}>0$ (positive sample) and 0 otherwise; $\sigma_{IoU}$ is the temporal Intersection over Union (IoU) between the predicted segment and the ground truth, serving as a weighting factor; $\mathcal{L}_{cls}$ is the classification loss, implemented as the focal loss Ross and Dollár (2017); and $\mathcal{L}_{reg}$ is the regression loss, implemented as the IoU loss Rezatofighi et al. (2019). The weighting factor $\sigma_{IoU}$ emphasizes predictions with higher temporal IoU, ensuring that higher-quality predictions contribute more significantly during training. Positive samples are determined using center sampling, where instants near the center of an action instance are labeled as positive, and others are considered negative.
Anomaly Detection Loss $\mathcal{L}_{AD}$ (MGFN Chen et al. (2023)):
@@
with $y$ as the video-level label ($y=1$ for anomaly, $y=0$ for normal) and $s^{i,j}$ as the computed anomaly score for frame $i$ in segment $j$. Following Sultani et al. (2018), it incorporate a temporal smoothness term $\mathcal{L}_{ts}$ and a sparsity term $\mathcal{L}_{sp}$:
@@
where $B$ is the batch size, $M$ denotes the feature magnitude of the corresponding segment, $l$ is an indicator function, and $D(\cdot,\cdot)$ is a distance function. Refer to Chen et al. (2023) for more details.
@@ Appendix C Additional Experiments
Num MHSA Heads VISPR HMDB51 UCF Crime
@@
Figure 3 follows Singh et al. (2022) to plot privacy-utility curve (NHV=0.6833) using PA-HMDB test set. Varying task weights leads to controllable trade-off curve.
@@
Figure 4: Qualitative reconstruction analysis using the VideoMAE decoder. We evaluate four input configurations: (a) reconstruction from non-anonymized full patch tokens, (b) reconstruction from patch tokens after applying our $f_{A}$, (c) reconstruction using repeated CLS tokens, and (d) reconstruction using anonymized CLS tokens. As shown, meaningful spatial structure is recoverable only when full patch tokens are provided; anonymized patches and CLS-only variants fail to reconstruct identifiable content.
Table 13: UCF101 retrieval results (query: test set, gallery: train set). Evaluation uses Recall@1, Recall@5, Recall@10, and mAP. Our anonymizer faithfully maintains retrieval performance.
Method R@1 R@5 R@10 mAP
Baseline (Raw Videos) 95.3 97.6 98.7 84.9
Ours 94.9 97.6 98.6 84.7
Video Retrieval: Table 13 shows UCF101 retrieval performance using an anonymizer trained solely on Kinetics400, demonstrating cross-dataset and cross-task generalization without any retrieval-specific supervision.
Reconstruction Attack Mitigation: Figure 4 illustrates our reconstruction test protocol, designed to evaluate whether our approach mitigates reconstruction-based attacks and if CLS embeddings contain sufficient information for pixel-space recovery. Using the official VideoMAE decoder, we compare four conditions: (a) reconstruction from the full set of patch tokens, which recovers the scene faithfully; (b) reconstruction after applying our $f_{A}$ to the patchwise tokens, which severely disrupts spatial content; (c) reconstruction when the decoder is retrained to operate on a repeated CLS token, which produces only faded scenes; and (d) reconstruction from our anonymized CLS embeddings, which similarly fails to yield meaningful structure. Each setting finetunes the decoder for 10 epochs on the UCF101 training set, ensuring fair comparison. Impressively, despite being trained to remove private attributes from CLS tokens, our method prevents reconstruction attacks from recovering the original video, even when provided patch-level tokens. We also confirm that CLS features lack the spatial detail necessary for reconstructing privacy-revealing video content.
@@ C.1 Comparison with Other Prior Methods
Previous work Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023) has already shown that the learnable anonymization techniques outperform methods such as downsampling, blurring, and blackening. Main Table 1 shows a comparison to these techniques using the I3D Carreira and Zisserman (2017) model. In Downsample-2x and Downsample-4x, the input frames have their resolution reduced by a factor of 2 ($112\times 112$) and 4 ($56\times 56$). In Blackening and Blurring, subjects are detected using an object detector to detect human subjects and obfuscated using the same methods as described in Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023). We see that none of these techniques achieve an acceptable level of anonymization, and almost all reduce utility more than our SPLAVU method, further demonstrating the capability of our framework.
Table 14 shows the results of our proposed method on PA-HMDB Wu et al. (2020) compared to the baseline model on raw data.
@@ C.2 Additional Experiments with Large Foundation Models
Due to the low compute cost and focus on maintaining the capabilities of powerful models, our SPLAVU framework is able to scale up to the largest video foundational models currently available. Table 15 demonstrates the high privacy-utility tradeoff achieved by our method using InternVideo-H Wang et al. (2022), and Main Paper Table 1 shows results using VideoMAEv2-G Wang et al. (2023). In these experiments, action recognition performance was exactly maintained, and private attribute prediction was dropped more than for the smaller models, with only a modest reduction in temporal action detection performance.
@@
Anonymization Method Model VISPR HMDB51 UCF101 THUMOS14
@@ C.2.1 Training Compute
One of the many benefits of our SPLAVU framework is its very low compute/training cost. Table 16 shows the overall count of trainable parameters for previous frameworks compared to our AAM. For VideoMAE-Base, our SPLAVU framework with the self-attention AAM has 88.7% less trainable parameters when compared to existing approaches. This difference is even greater when scaling to larger models. Training less parameters can reduce the tendency to overfit on the proxy task and allow for learning an effective anonymization on limited training data (see Main Paper Table 4). Also, in federated learning, these parameters are communicated between the server and clients, so the reduced learnable parameters are useful in efficient and privacy-preserving federated learning Zhao et al. (2023); Yu et al. (2022).
@@ C.2.2 Precomputing Feature Embeddings
Training Data PAP AR TAD AD Training
@@ Appendix D Training Algorithm
Algorithm 1 formalizes the SPLAVU workflow notation. We consider anonymizer $f_{A}$ and task heads $f_{T_{AR}}$, $f_{T_{TAD}}$, and $f_{T_{AD}}$ for the anonymization training and $f_{AR}$, $f_{TAD}$, and $f_{wsad}$ for downstream tasks. In order, these models are parameterized by $\theta_{A}$, $\theta_{{T}_{AR}}$, $\theta_{{T}_{TAD}}$, $\theta_{{T}_{AD}}$, $\theta_{AR}$, $\theta_{TAD}$, and $\theta_{wsad}$. $\mathbb{D}_{AR}$, $\mathbb{D}_{TAD}$, and $\mathbb{D}_{wsad}$ are all used in the proxy anonymization process, then also for the downstream task evaluation. The downstream $\mathbb{D}_{AR}$ may be the same or different from during the anonymization process.
Anonymization Training
Inputs:
Datasets: $\mathbb{D}_{AR}$, $\mathbb{D}_{TAD}$, $\mathbb{D}_{wsad}$
# of Epochs: anon_epochsanon\_epochs
Learning Rates: $\alpha_{A}$, $\alpha_{AR}$, $\alpha_{TAD}$, $\alpha_{AD}$
Hyperparameters: $\omega_{A}$, $\omega_{T}$, $\omega_{B}$, $\omega_{LC}$, $\omega_{AR}$, $\omega_{TAD}$, $\omega_{AD}$
Output: $\theta_{A}$, $\theta_{{T}_{AR}}$, $\theta_{{T}_{TAD}}$, $\theta_{{T}_{AD}}$
Model Initialization:
Initialize $f_{E}$ with Kinetics400 weights Carreira and Zisserman (2017);
Initialize $\theta_{A}\leftarrow\theta_{A}-\alpha_{A}\nabla_{\theta_{A}}(\mathcal{L}_{L1}(\theta_{A}))$
Multitask Anonymization Training:
for $e_{0}\leftarrow 1$ to anon_epochsanon\_epochs do
 $\theta_{A}\leftarrow\theta_{A}-\alpha_{A}\nabla_{\theta_{A}}(\omega_{LC}\mathcal{L}_{LC}(\theta_{A})+\omega_{T}\mathcal{L}_{T^{*}}(\theta_{A},\theta_{T_{AR}},\theta_{T_{TAD}},\theta_{T_{AD}})-\omega_{B}L_{B}(\theta_{A}))$ $\theta_{T_{AR}}\leftarrow\theta_{T_{AR}}-\alpha_{AR}\nabla_{\theta_{T_{AR}}}(\mathcal{L}_{AR}(\theta_{T_{AR}},\theta_{A})),$ $\theta_{T_{TAD}}\leftarrow\theta_{T_{TAD}}-\alpha_{TAD}\nabla_{\theta_{T_{TAD}}}(\mathcal{L}_{TAD}(\theta_{T_{TAD}},\theta_{A})),$ $\theta_{T_{AD}}\leftarrow\theta_{T_{AD}}-\alpha_{AD}\nabla_{\theta_{T_{AD}}}(\mathcal{L}_{AD}(\theta_{T_{AD}},\theta_{A})),$
end for
 Downstream Tasks Evaluation
Inputs:
Datasets: $\mathbb{D}_{AR}$, $\mathbb{D}_{AD}$, $\mathbb{D}_{TAD}$
# of Epochs: reco_epochsreco\_epochs, anomaly_epochsanomaly\_epochs, tad_epochstad\_epochs
Learning Rates: $\alpha_{AR}$, $\alpha_{wsad}$, $\alpha_{TAD}$
Output: $\theta_{AR}$, $\theta_{wsad}$, $\theta_{TAD}$
 Privacy-Preserved Action Recognition Training:
for $e_{0}\leftarrow 1$ to reco_epochsreco\_epochs do
 $\theta_{AR}\leftarrow\theta_{AR}-\alpha_{AR}\nabla_{\theta_{AR}}(\mathcal{L}_{T}(\theta_{AR},\theta_{A})),$
end for
 Feature Extraction on $\mathbb{D}_{AD}$:
$\mathbb{F}_{AD}=\{\,f_{A}(f_{E}(X^{(i))}))\;|\;\forall X^{(i)}\in\mathbb{D}_{AD}\,\}$
@@
 Privacy-Preserved Weakly-Supervised Anomaly Detection (WSAD) Training:
for $e_{0}\leftarrow 1$ to anomaly_epochsanomaly\_epochs do
 $\theta_{wsad}\leftarrow\theta_{wsad}-\alpha_{wsad}\nabla_{\theta_{wsad}}(L_{wsad}(\theta_{wsad},\mathbb{F}_{AD}))$
end for
 Feature Extraction on $\mathbb{D}_{TAD}$:
$\mathbb{F}_{TAD}=\{\,f_{A}(f_{E}(X^{(i)}))\;|\;\forall X^{(i)}\in\mathbb{D}_{TAD}\,\}$
 Privacy-Preserved Temporal Action Detection (TAD) Training:
for $e_{0}\leftarrow 1$ to tad_epochstad\_epochs do
 $\theta_{TAD}\leftarrow\theta_{TAD}-\alpha_{TAD}\nabla_{\theta_{TAD}}(L_{TAD}(\theta_{TAD},\mathbb{F}_{AD}))$
end for